# High resolution quantum enhanced phase imaging of cells

### v1

@@ Abstract
Recovering both amplitude and phase information from a system is a fundamental goal of optical imaging. At the same time, it is crucial to use a low photon dose to avoid altering the system, particularly when dealing with biological samples. Quantum imaging offers a powerful approach for extracting more information per photon than classical techniques, which are ultimately limited by shot-noise. However, the trade-off between quantum noise reduction and spatial resolution has been considered a major drawback to the application of quantum techniques to small cellular and sub-cellular structures, where they could offer the most significant benefits. In this work, we overcome this limitation by demonstrating a resolution-independent quantum advantage. We achieve high-resolution phase imaging limited only by the numerical aperture, while simultaneously attaining quantum noise reduction. This enables, for the first time, sub-shot-noise quantitative phase imaging of biological cells. Unlike other quantum imaging approaches, our method operates in a quasi-single-shot wide-field mode, retrieves both phase and amplitude information, and does not rely on interferometric measurements, making it intrinsically fast and stable. These results pave the way for the immediate application of sub-shot-noise imaging in biology.
@@ I Introduction
In biological imaging, many specimens, such as live cells and tissues, are nearly transparent under conventional light microscopy, which makes high-contrast imaging challenging without artificial labeling. Quantitative phase imaging (QPI) addresses this limitation by measuring optical path length differences, providing intrinsic contrast based on sample thickness and refractive index variations Park et al. (2018).
@@
In general, extracting meaningful information from both the amplitude and phase of an object depends critically on two key figures of merit: resolution and sensitivity. Resolution defines the smallest distinguishable details in an image and is influenced by factors such as the optical system, the wavelength of light, and detector properties. Sensitivity, on the other hand, refers to the ability to detect weak signals, which is crucial for imaging faint biological structures. Sensitivity is often limited by photon quantum noise (shot-noise) and becomes a bottleneck when increasing the illumination level, which may induce phototoxicity or interfere with the biological processes under study. Often, improving resolution may be detrimental to sensitivity and vice versa, while it is highly desirable to maintain a balance between the two.
@@
Quantum sensing Pirandola et al. (2018); Degen et al. (2017); Petrini et al. (2020); Lee et al. (2021); Polino et al. (2020) and imaging Defienne et al. (2024); Berchera and Degiovanni (2019); Genovese (2016) offer a solution to these limitations, with applications spanning from fundamental physics Aasi et al. (2013); Pradyumna et al. (2020) to biological measurements Taylor et al. (2014); Casacio et al. (2021); Petrini et al. (2022), target detection Tan et al. (2008); Lopaeva et al. (2013); Zhang et al. (2015); Gallego Torromé and Barzanjeh (2024), optical memory readout Ortolano et al. (2021a, b), and pattern recognition Ortolano et al. (2023a).
In quantum imaging, single-photon emitters have enabled super-resolution fluorescence microscopy Gatto Monticone et al. (2014); Schwartz and Oron (2012); Tenne et al. (2019), while label-free resolution enhancement has been achieved using position-momentum correlated photon pairs Toninelli et al. (2019); Defienne et al. (2022); Zhang et al. (2024) and coincidence measurements, leveraging multi-photon quantum coherence D’Angelo et al. (2017); Unternährer et al. (2018), which underpins quantum lithography D’Angelo et al. (2001). Due to the intrinsic spatial and temporal correlations of photon pairs, uncorrelated external background Lopaeva et al. (2013); Defienne et al. (2019); Gregory et al. (2020); Zhang et al. (2024) can also be rejected.
Quantum techniques also enable new phase imaging schemes: spatial and polarization entanglement have been used for quantum holography Töpfer et al. (2022); Black et al. (2023); Defienne et al. (2021), imaging with undetected photons Lemos et al. (2014); Kviatkovsky et al. (2020); Vallés et al. (2018), and S(1,1) nonlinear interferometry Machado et al. (2020), which enables separation between the probing and detection wavelengths. Alternatively, quantum methods such as Hong-Ou-Mandel interferometry allow depth profiling via photon indistinguishability Devaux et al. (2020); Ndagano et al. (2022). However, these approaches do not enhance sensitivity beyond the shot-noise limit of their classical counterparts.
Phase supersensitivity can be achieved by exploiting $N$-photon entangled NOON states, which exhibit a de Broglie wavelength of $\lambda/N$ Giovannetti et al. (2011). While proof-of-principle experiments have been conducted using 2-photon entanglement—initially with raster-scanning systems Ono et al. (2013); Israel et al. (2014) and more recently in wide-field modalities Camphausen et al. (2021)—practical implementations remain challenging due to the difficulties in generating NOON states with $N>2$ and their susceptibility to optical losses. In general, lateral spatial resolution remains constrained by the biphoton correlation width, which is tied to maximizing the joint two-photon detection probability.
@@
A different approach, called sub-shot-noise quantum imaging (SSNQI) Brida et al. (2010); Samantaray et al. (2017); Sabines-Chesterking et al. (2019); Ruo-Berchera et al. (2020), surpasses the sensitivity limit in weak object transmission measurements, ideally by a factor $\sqrt{1-\tau}$ Monras and Paris (2007); Adesso et al. (2009); Nair (2018); Losero et al. (2018); Wang et al. (2024), where $0\leq\tau\leq 1$ is the sample transmittance. SSNQI exploits non-classical intensity correlations between two spatially separated beams Jakeman and Rarity (1986): one interacts with the object forming a noisy classical image, while the other serves as a quantum noise reference, which is subtracted from the object image. Recent advances have extended SSNQI to non-interferometric quantum-enhanced phase imaging (NIQPI) Ortolano et al. (2023b), where quantitative phase information is retrieved from intensity patterns using the transport-of-intensity equation (TIE) Teague (1983); Paganin and Nugent (1998); Zuo et al. (2020). A similar non-interferometric approach was recently employed for biphoton state reconstruction Dehghan et al. (2024).
The strength of SSNQI and NIQPI lies in their favorable operating conditions: they do not require phase stability as in interferometric techniques, operate rapidly with quasi single-shot first-order intensity measurements, and are completely scanning-free.
However, earlier demonstrations were limited by a significant trade-off between noise reduction and spatial resolution. Specifically, strong shot-noise suppression was achieved by maximizing the collection of correlated photon pairs using detection areas larger than the photon-pair transverse spatial correlation area. While this optimization enhanced noise reduction, it inevitably compromised spatial resolution, effectively acting as a low-pass filter and blurring fine details.
Here, we show that, contrary to prior assumptions, NIQPI’s resolution is not fundamentally constrained by such averaging. This breakthrough arises from the specific and favorable interplay between the features of SSNQI and the TIE-based reconstruction algorithm, enabling excellent quantum noise reduction while preserving fine details, with resolution virtually limited only by the numerical aperture of the imaging system.
Importantly, we demonstrate this quantum advantage in samples containing both phase and amplitude structures, achieving noise reduction exceeding 30%.
Finally, we report wide-field quantum-enhanced quantitative phase imaging of thin biological samples, achieving cellular and sub-cellular resolution. This opens a concrete perspective for immediate application and the potential development of commercial quantum imaging prototypes.
@@
Figure 1: High resolution quantum phase imaging technique (a) Schematic of the non-interferometric quantum phase imaging setup Ortolano et al. (2023b). Signal and idler beams generated by spontaneous parametric down conversion (SPDC) propagate through an $f$-$f$ optical system, producing correlated intensity patterns. A phase-amplitude object, placed near the far field of the source, interacts only with the signal beam. Phase information is retrieved from intensity measurements placing the object at the focal and at the defocused planes (±$\pm dz$). (b) Engineered test object consisting of distinct phase and transmittance masks. The phase mask consists of the superposition of a pure phase structure featuring a $\pi$-shaped symbol and a $\varnothing$-shaped symbol. The $\varnothing$ symbol additionally features a transmittance of approximately 0.94. (c) Experimental and simulated quantum advantage in phase estimation using the transmittance-optimized correction factor $k^{(\tau)}_{\text{opt}}$ Ortolano et al. (2023b) and the newly proposed factor for TIE-based phase retrieval $k^{(TIE)}_{\text{opt}}$. The advantage is quantified as the ratio $C(\text{Quant.})/C(\text{Clas.})$. ), where $C(\text{Quant.})$. ) ($C(\text{Clas.}))$. ) ) are the correlation coefficients between the quantum (classical) phase images and the reference one. The ratio is plotted as a function of the defocus distance $dz$ and the resolution factor $\mathcal{D}=L_{\text{det}}/(M\cdot l_{\text{CFF}})$, with $L_{\text{det}}$ the linear size of the integration area, $M$ the magnification of the imaging system, and $l_{\text{CFF}}$ the correlation length. The different acquired values of $\mathcal{D}$ are obtained by varying $L_{\text{det}}$. Red and blue dots represent experimental data using $k_{\tau}$ and $k_{\text{TIE}}$, respectively, with shaded regions showing theoretical predictions. Green highlights indicate areas of optimization gain.
@@
The SSNQI and NIQPI protocols exploit the scheme depicted in Fig. 1a Ortolano et al. (2023b). In a type-II nonlinear crystal pumped by a Gaussian coherent laser beam, spontaneous parametric down-conversion (SPDC) generates signal ($s$) and idler ($i$) beams, which become spatially separated at the crystal’s far-field (CFF), obtained at the focus of a lens in an $f$-$f$ configuration. Signal and idler photons are always generated in pairs along correlated and symmetric directions with respect to the pump propagation axis. This angular correlation translates into a transverse spatial correlation in the CFF, characterized by a correlation length $l_{\mathrm{CFF}}$ ($l_{\mathrm{CFF}}\approx 5\,\mu\mathrm{m}$ in our setup, approximately symmetric along both transverse directions). The corresponding correlation area, $l_{\mathrm{CFF}}^{2}$ in the other beam, is approximately four orders of magnitude smaller than the overall spot size of each beam. In the low-gain regime of SPDC, the intensity patterns of the $s$- and $i$-beams are each characterized by shot-noise spatial fluctuations with a white spatial spectrum, i.e., the shot-noise is equally present at all spatial frequencies. Furthermore, the shot-noise fluctuations of the two beams are almost identically correlated (at spatial frequencies larger than $l_{\text{CFF}}$) across the entire spot size, a feature impossible to reproduce with classical light sources.
The signal beam probes the object placed near the CFF plane. An imaging system projects the CFF onto a low-noise, high quantum-efficiency charge-coupled device (CCD) camera with magnification $M$, yielding photon counts $N_{s}(\mathbf{x},dz)$ in the pixel with transverse coordinate $\mathbf{x}$, where $dz$ denotes the distance of the object from the CFF along the propagation axis. Hereinafter, we will indicate by $N^{\prime}_{s}$, with (), the intensity measured with the object and $N_{s}$ the one without the object. Simultaneously, the idler beam is directed onto a separate region of the detector, producing a photon-count distribution $N_{i}(\mathbf{x})$, which is used to subtract the quantum-correlated shot-noise from the measured intensity in the signal arm. These intensity correlations can be directly applied in the SSNQI protocol. In particular, the transmission of an object placed in the CFF plane, $\tau(\mathbf{x})$, is estimated as Moreau et al. (2017); Ruo-Berchera et al. (2020):
@@
NIQPI, presented in Ortolano et al. (2023b), successfully applied SSNQI to image pure-phase samples using the TIE. While a pure-phase object produces no intensity contrast when precisely located in the focal plane of the imaging system, a slight displacement from this plane induces intensity variations due to propagation effects stemming from spatial phase gradients. The TIE method exploits these propagation-induced intensity changes to quantitatively retrieve the phase profile by solving the following equation:
@@
which links the gradient of the transverse phase, denoted by $\phi(\mathbf{x},0)$, on the right-hand side to the axial derivative of the intensity on the left-hand side. The TIE formulation holds under the paraxial approximation and remains valid even for partially coherent illumination Paganin and Nugent (1998).
In practice, the derivative in Eq. (2) is approximated by a finite difference using two intensity measurements taken at planes symmetrically displaced by ±$\pm dz$:
@@
Ideally, choosing an infinitesimally small displacement ($dz\rightarrow 0$) would yield the most accurate approximation of the derivative. However, smaller values of $dz$ significantly reduce the strength of phase-induced intensity variations, making the signal indistinguishable from electronic and fundamental shot-noise. Conversely, larger displacements enhance noise resilience but reduce the accuracy of the finite-difference approximation, leading to out-of-focus phase reconstructions.
@@
Figure 2: Experimental reconstruction of transmittance and phase object (a) Transmittance estimations for classical and quantum single-frame images, processed with a factor $\mathcal{D}=3.8$. The object’s area occupies $220\times 220~{}\text{pix}^{2}$ with a mean number of photon per pixel of 600. (b) Phase reconstruction at various defocus distances, $dz$, comparing classical and quantum single-frame approaches. The phase estimations at the smallest $dz$ are: classical $\pi=(-0.260\pm 0.030$) rad, quantum $\pi=(-0.259\pm 0.022)$ rad (theoretical $\pi$ = $(-0.226\pm 0.006)$ rad); classical $\varnothing$ $=(0.365\pm 0.020)$ rad, quantum $\varnothing$ $=(0.364\pm 0.015)$ rad (theoretical $\varnothing$ = $(0.345\pm 0.014)$ rad). Experimental uncertainties are obtained as mean errors calculated over $10^{3}$ frames.
@@
Thus, an optimal value of $dz$ must balance the accuracy of the derivative approximation against robustness to noise, and its choice critically depends on both the noise level and the specific phase profile of the sample. In this context, reducing shot-noise is essential for accurately approximating the intensity derivative via finite differences, thereby improving both the quantitative accuracy and qualitative clarity of the retrieved phase images. Details about the solution of Eq. (2) are provided in the M&M section.
@@
In NIQPI, the quantum noise reduction is performed analogously to the SSNQI approach, substituting intensity terms in Eq. (2) with quantum-corrected intensities Ortolano et al. (2023b):
$I(\mathbf{x},z)\mapsto N^{\prime}_{s}(\mathbf{x},z)-k^{(\text{TIE})}_{\text{opt}}\,\delta N_{i}(\mathbf{x},0),$ (4)
where the number of photon counts, $N$, is considered in place of the intensities $I$ since the two quantities are proportional, given a fixed detection time and area.
@@
While previous experiments focused exclusively on pure-phase samples, extending the approach to realistic biological applications introduces additional complexities. Biological specimens typically feature regions dominated by phase variations with negligible absorption, alongside areas where phase and absorption effects coexist. Accurately retrieving and distinguishing these two types of information is essential for a comprehensive characterization of both the structural and functional features of biological samples.
@@
To address this general scenario, we employed an engineered test object composed of distinct phase and transmittance masks (Fig. 1b). The phase mask consists of the superposition of a pure phase structure featuring a $\pi$-shaped symbol and a $\varnothing$-shaped symbol (more information is provided in the Supplementary Document). The respective phase shifts are $(\pi=-0.226\pm 0.006)$ rad and $(\varnothing=0.345\pm 0.014)$ rad relative to the background @810 nm, the degenerate wavelength of the SPDC process. Additionally, the $\varnothing$-shaped symbol exhibits a transmittance of approximately 0.94 relative to the flat of the Silica.
@@
The TIE method inherently addresses these challenges by decoupling phase-induced intensity variations from those arising due to the non-uniform transmittance Zuo et al. (2020), thanks to the use of the intensity measurements acquired at symmetrically defocused planes, $I(\mathbf{x},\pm dz)$, as further described in the M&M section.
@@
The noise reduction factor (NRF), defined as $\text{NRF}=~{}\frac{\langle\delta^{2}(N_{s}-N_{i})\rangle}{\langle N_{s}+N_{i}\rangle},$ is a well-established quantifier of non-classical correlations and shot-noise suppression, comparing the residual fluctuations in the difference of photon counts (numerator) to the shot-noise level expected for classical states (denominator) Bondani et al. (2007); Brida et al. (2009); Jedrkiewicz et al. (2004); Blanchet et al. (2008); Agafonov et al. (2010). For classical light, the NRF satisfies $\text{NRF}\geq~{}1$, while quantum-correlated states can reach values in the sub-shot-noise regime, $0\leq\text{NRF}<1$.
In the case of two correlated regions of a detector collecting SPDC photons, the NRF can be expressed as $\text{NRF}=1-\eta_{0}\eta_{c}(\mathcal{D}),$ where $\eta_{0}$ is the single-channel detection efficiency, with the assumption of a balanced setup, $\eta_{s}=\eta_{i}=\eta_{0}$, and $\eta_{c}(\mathcal{D})$ represents the collection efficiency of correlated photon pairs Samantaray et al. (2017); Ruo-Berchera et al. (2020). The quantity $\eta_{c}(\mathcal{D})$ is a increasing function of the ratio $\mathcal{D}=\frac{L_{\text{det}}}{M\cdot l_{\text{CFF}}},$ with $L_{\text{det}}$ the linear size of the detector’s integration area, $M$ the system magnification, and $l_{\text{CFF}}$ the correlation length in the crystal far-field. Further details about the estimation of $\eta_{0}$ and $\eta_{c}(\mathcal{D})$ are provided in the M&M section. When $\mathcal{D}\gg 1$, most of the correlated photon pairs are collected by two symmetric detection areas in the signal and idler beams, maximizing $\eta_{c}(\mathcal{D})\approx 1$; conversely, for $\mathcal{D}\lesssim 1$, only a fraction of the pairs are detected.
It is important to note that, while the spatial resolution $r_{\mathrm{CFF}}$ of the signal beam alone is inherently limited by the numerical aperture of the imaging system to $r_{\mathrm{CFF}}\approx 1.5\,\mu\mathrm{m}$ in our setup, noise subtraction in the intensity image is effective only on spatial scales larger than $l_{\mathrm{CFF}}$, that is for $\mathcal{D}\gtrsim 1$. This introduces a fundamental trade-off between spatial resolution and quantum advantage in SSNQI for amplitude objects, based on the estimator in Eq. (1). This dependency has been extensively characterized in previous works Samantaray et al. (2017); Meda et al. (2017); Ruo-Berchera et al. (2020). Under the assumptions of quasi-Poissonian photon number statistics in each arm and high object transmittance $\tau\approx 1$, the optimization factor $k^{(\tau)}_{\text{opt}}$ can be approximated as $k^{(\tau)}_{\text{opt}}(\mathcal{D})\approx 1-\text{NRF}=\eta_{0}\eta_{c}(\mathcal{D}),$ indicating the dependence on the resolution parameter $\mathcal{D}$. In practice, $k^{(\tau)}_{\text{opt}}$ is calibrated through a preliminary characterization of the experimental NRF as a function of $\mathcal{D}$ (see M&M).

### v2

@@ Abstract
Recovering both amplitude and phase information from a system is a fundamental goal of optical imaging. At the same time, it is crucial to operate at low photon doses to avoid altering the sample, particularly in biological applications. Quantum imaging provides a powerful route to extract more information per photon than classical techniques, which are ultimately limited by shot-noise. However, the trade-off between quantum noise reduction and spatial resolution has long been regarded as a major obstacle to the application of quantum techniques to small cellular and sub-cellular structures, where they could offer the greatest benefits. Here, we overcome this limitation by demonstrating sub-shot-noise quantitative phase imaging of biological cells based on the transport-of-intensity equation, enabling high-fidelity, label-free imaging of key cellular and sub-cellular features. We achieve high-resolution phase imaging limited only by the numerical aperture, while simultaneously obtaining a resolution-independent quantum advantage. Unlike other quantum imaging approaches, our method operates in a quasi-single-shot, wide-field configuration, retrieves both phase and amplitude information, and does not rely on interferometric measurements, making it intrinsically fast and stable. These results pave the way for the immediate application of sub-shot-noise imaging in biological microscopy.
@@ I Introduction
In biological imaging, many specimens, such as live cells and tissues, are nearly transparent under conventional light microscopy, making high-contrast imaging challenging without artificial labeling. Quantitative phase imaging (QPI) addresses this limitation by measuring optical path length differences, providing intrinsic contrast based on sample thickness and refractive index variations [48].
@@
In general, the extraction of meaningful information from both the amplitude and phase of an object critically depends on two key figures of merit: resolution and sensitivity. Resolution determines the smallest distinguishable detail in an image and is influenced by factors such as the optical system, the wavelength of light, and detector properties. Sensitivity, by contrast, refers to the ability to detect weak signals and is crucial for imaging faint biological structures. In optical imaging, sensitivity is often limited by photon quantum noise (shot-noise) and becomes a bottleneck when increasing the illumination level may induce phototoxicity or interfere with the biological processes under investigation [59, 12]. Improving resolution is often detrimental to sensitivity and vice versa, whereas maintaining an optimal balance between the two remains highly desirable.
@@
Quantum sensing [51, 19, 49, 31, 52] and imaging [15, 4, 24] offer a solution to these limitations, with applications spanning from fundamental physics [1, 53] to biological measurements [60, 11, 50], target detection [58, 33, 70, 22], optical memory readout [43, 42], and pattern recognition [44].
In quantum imaging, single-photon emitters have enabled super-resolution fluorescence microscopy [23, 57, 62]. Entangled photon pairs have been used for label-free spatial resolution enhancement [63, 16, 69, 14, 65], quantum lithography [13], external background rejection [33, 18, 26, 69]. Quantum interferometric sensing approaches includes holography [64, 5, 17], imaging with undetected photons [32, 30, 66], S(1,1) nonlinear interferometry [35] and Hong-Ou-Mandel depth profiling [21, 40]. However, these approaches do not enhance sensitivity beyond the shot-noise limit of their classical counterparts. Phase sensitivity beyond the standard quantum limit can be achieved by exploiting $N$-photon entangled NOON states, which exhibit a de Broglie wavelength of $\lambda/N$ [25]. Despite proof-of-principle experiments [41, 27, 10], practical implementations remain challenging due to the difficulties in generating NOON states with $N>2$ and their susceptibility to optical losses and decoherence.
@@
A more practical approach is provided by sub-shot-noise quantum imaging (SSNQI) [8, 56, 55, 54], which enables sensitivity beyond the classical limit in weak object transmission measurements [37, 2, 39, 34, 68]. SSNQI exploits non-classical intensity correlations between two spatially separated beams [28]: one beam interacts with the object forming a noisy classical image, while the other serves as a quantum noise reference, which is subtracted from the object image. Recent advances have extended SSNQI to non-interferometric quantum-enhanced phase imaging (NIQPI) [45], where quantitative phase information is retrieved from intensity patterns using the transport-of-intensity equation (TIE) [61, 47, 72]. A similar non-interferometric approach was recently employed for biphoton state reconstruction [20]. The strength of SSNQI and NIQPI lies in their favorable operating conditions: they do not require phase stability as in interferometric techniques, operate rapidly with quasi-single-shot first-order intensity measurements, and are completely scanning-free. However, earlier demonstrations  [8, 56, 55, 45] were limited by a significant trade-off between noise reduction and the lateral spatial resolution, which remains constrained by the biphoton correlation width at the object plane.
In this work, we introduce a high-resolution NIQPI approach that overcomes the resolution–sensitivity trade-off, achieving the first demonstration of quantitative sub-shot-noise phase imaging of biological cells at high spatial resolution. We identify and address the key mechanism underlying this advance: a previously unexplored interplay between noise propagation in TIE-based phase reconstruction and the specific properties of quantum noise subtraction. We provide a rigorous characterization of the experimental setup in terms of phase reconstruction fidelity, spatial resolution, and quantum advantage, using a pre-characterized nanofabricated sample featuring both phase and amplitude structures. We demonstrate a quantum advantage of approximately 30% in reconstruction fidelity at a fixed illumination level. More importantly, we establish the immediate applicability of this technique by imaging unstained sea urchin ova, revealing cellular and sub-cellular features with a quantum-enhanced signal-to-noise ratio. Overall, this work bridges the gap between proof-of-concept quantum metrology and practical biological imaging, opening a concrete pathway toward the development of deployable quantum imaging technologies.
@@
Figure 1: High resolution quantum phase imaging technique (a) Schematic of the non-interferometric quantum phase imaging setup [45]. Signal and idler beams generated by spontaneous parametric down conversion (SPDC) propagate through an $f$-$f$ optical system, producing correlated intensity patterns. A phase-amplitude object, placed near the far field of the source, interacts only with the signal beam. Phase information is retrieved from intensity measurements placing the object at the focal and at the defocused planes ($\pm dz$). (b) Engineered test object consisting of distinct phase and transmittance masks. The phase mask consists of the superposition of a pure phase structure featuring a $\pi$-shaped symbol and a $\varnothing$-shaped symbol. The $\varnothing$ symbol additionally features a transmittance of approximately 0.93. The phase-amplitude object has lateral dimensions of approximately $348\,\mu\mathrm{m}\times 350\,\mu\mathrm{m}$, with thicknesses of about $60$ nm for the $\pi$ structure and $40$ nm for the $\varnothing$ symbol. (c) The four panels show the experimental and simulated quantum advantage in phase estimation using the transmittance-optimized correction factor $k^{(\tau)}_{\text{opt}}$ [45] and the newly proposed factor for TIE-based phase retrieval $k^{(TIE)}_{\text{opt}}$ for four different defocusing distances $dz$. The advantage is quantified as the ratio $C(\text{Quant.})/C(\text{Clas.})$, where $\text{Quant.}) (C(C$las.))C($\text${Clas.})) are the correlation coefficients between the quantum (classical) phase images and the reference one. The reported mean values and the associated uncertainty bars, corresponding to one standard uncertainty on the mean, are evaluated over an ensemble of approximately $10^{3}$ images. The ratios, for the different distances $dz$, are plotted as a function of the phase spatial resolution. Red and blue dots represent experimental data using $k^{(\tau)}_{\text{opt}}$ and $k^{(TIE)}_{\text{opt}}$, respectively, with shaded regions showing theoretical predictions. Green highlights indicate areas of optimization gain.
@@
The SSNQI and NIQPI protocols exploit the scheme depicted in Fig. 1a [45]. In a type-II nonlinear crystal pumped by a Gaussian coherent laser beam, spontaneous parametric down-conversion (SPDC) generates signal ($s$) and idler ($i$) beams, which become spatially separated in the crystal’s far-field (CFF), obtained at the focus of a lens in an $f$-$f$ configuration. Signal and idler photons are always generated in pairs along correlated and anti-symmetric directions leading to anti-correlated positions in the CFF. The spatial correlation in the CFF is characterized by a correlation length $l_{\mathrm{CFF}}$ ($l_{\mathrm{CFF}}\approx 5\,\mu\mathrm{m}$ in our setup [56], approximately symmetric along both transverse directions, see M&M for further details). The corresponding correlation area, $l_{\mathrm{CFF}}^{2}$, is approximately four orders of magnitude smaller than the overall spot size of each beam. In the low-gain regime of SPDC, the intensity patterns of the $s$- and $i$-beams are each characterized by shot-noise spatial fluctuations with a white spatial spectrum, i.e., the shot-noise is equally present at all spatial frequencies. Furthermore, the shot-noise fluctuations of the two beams are almost identically correlated (at spatial areas larger than $l_{\text{CFF}}$) across the entire spot size, a feature impossible to reproduce with classical light sources.
The signal beam probes the object placed near the CFF plane. An imaging system projects the CFF onto a low-noise, high quantum-efficiency charge-coupled device (CCD) camera with magnification $M$, yielding photon counts $N_{s}(\mathbf{x},dz)$ in the pixel at the transverse coordinate $\mathbf{x}$, where $dz$ denotes the distance of the object from the CFF along the propagation axis. Hereinafter, we will indicate by $N^{\prime}_{s}$, with (), the intensity measured with the object and $N_{s}$ the one without the object. Simultaneously, the idler beam is directed onto a separate region of the detector, producing a photon-count distribution $N_{i}(\mathbf{x})$, which is used to subtract the quantum-correlated shot-noise from the measured intensity in the signal arm. These intensity correlations can be directly applied in the SSNQI protocol. In particular, the transmission of an object placed in the CFF plane, $\tau(\mathbf{x})$, is estimated as [38, 54]:
@@
NIQPI, presented in [45], successfully applied SSNQI to image pure-phase samples using the TIE. While a pure-phase object produces no intensity contrast when precisely located in the focal plane of the imaging system, a slight displacement from this plane induces intensity variations due to propagation effects stemming from spatial phase gradients. The TIE method exploits these propagation-induced intensity changes to quantitatively retrieve the phase profile by solving the following equation:
@@
which links the gradient of the transverse phase, denoted by $\phi(\mathbf{x},0)$, on the right-hand side to the axial derivative of the intensity on the left-hand side. The TIE formulation holds under the paraxial approximation and remains valid even for partially coherent illumination [47].
In practice, the derivative in Eq. (2) is approximated by a finite difference using two intensity measurements taken at planes symmetrically displaced by $\pm dz$:
@@
In this sense, the technique is ‘quasi-single-shot’ since only three full-field intensity measurements performed at each plane (dz,0,dz-dz,0,dz) are required. Ideally, choosing an infinitesimally small displacement ($dz\rightarrow 0$) would yield the most accurate approximation of the derivative. However, smaller values of $dz$ significantly reduce the strength of phase-induced intensity variations, making the signal indistinguishable from electronic and fundamental shot-noise. Conversely, larger displacements enhance noise resilience but reduce the accuracy of the finite-difference approximation, leading to out-of-focus phase reconstructions.
@@
Figure 2: Experimental reconstruction of transmittance and phase object (a) Transmittance estimations for classical and quantum single-frame images, processed with a factor $\mathcal{D}=3.8$ obtained by averaging over 12 pixels. The object’s area occupies $220\times 220~\text{pix}^{2}$ with a mean number of photon per pixel of 600. (b) Phase reconstruction at various defocus distances, dzdz, comparing classical and quantum single-frame approaches. The phase estimations at the smallest dzdz are: classical $\pi=(-0.260\pm 0.030$) rad, quantum $\pi=(-0.259\pm 0.022)$ rad (theoretical $\pi$ = $(-0.226\pm 0.006)$ rad); classical $\varnothing$ $(0.37\pm 0.02)$ rad, quantum $\varnothing$ $(0.36\pm 0.02)$ rad (theoretical $\varnothing$ = $(0.35\pm 0.01)$ rad). Experimental uncertainties are obtained as mean errors calculated over $10^{3}$ frames.
@@
Therefore, an optimal value of $dz$ must balance the accuracy of the derivative approximation against robustness to noise, and its choice critically depends on both the noise level and the specific phase profile of the sample. In this context, reducing shot-noise is essential for accurately approximating the intensity derivative via finite differences, thereby improving both the quantitative accuracy and qualitative clarity of the retrieved phase images. Details about the solution of Eq. (2) are provided in the M&M section.
@@
In NIQPI, the quantum noise reduction is performed analogously to the SSNQI approach, replacing intensity terms in Eq. (2) with quantum-corrected intensities [45]:
$I(\mathbf{x},z)\mapsto N^{\prime}_{s}(\mathbf{x},z)-k^{\text{(TIE)}}_{\text{opt}}\,\delta N_{i}(\mathbf{x},0),$ (4)
where the number of photon counts, $N$, is considered in place of the intensities $I$ since the two quantities are proportional, given a fixed detection time and area. The classical reference is obtained by the substitution $I(\mathbf{x},z)\mapsto N^{\prime}_{s}(\mathbf{x},z)$ without applying any noise subtraction, thus ensuring a fair comparison at equal photon dose. Note that, in the low-gain regime, each SPDC beam exhibits Poissonian spatial photon-number fluctuations, so this classical reconstruction represents the optimal classical benchmark. A description of the measured statistical properties is reported in the M&M section.
@@
While previous experiment focused exclusively on pure-phase samples [45], extending the approach to realistic biological applications introduces additional complexities. Biological specimens typically feature regions dominated by phase variations with negligible absorption, alongside areas where phase and absorption effects coexist. Accurately retrieving and distinguishing these two types of information is essential for a comprehensive characterization of both the structural and functional features of biological samples.
@@
To address this general scenario, we employed an engineered test object composed of distinct phase and transmittance masks (Fig. 1b). The phase mask consists of the superposition of a pure phase structure featuring a $\pi$-shaped symbol and a $\varnothing$-shaped symbol (more information is provided in the Supplementary Document). The respective phase shifts are $(\pi=-0.226\pm 0.006)$ rad and $(\varnothing=0.35\pm 0.01)$ rad relative to the background @810 nm, the degenerate wavelength of the SPDC process. Additionally, the $\varnothing$-shaped symbol exhibits a transmittance of $(0.93\pm 0.01)$ relative to the flat silica region.
@@
The TIE method inherently addresses these challenges by decoupling phase-induced intensity variations from those arising due to the non-uniform transmittance [72], thanks to the use of the intensity measurements acquired at symmetrically defocused planes, $I(\mathbf{x},\pm dz)$, as further described in the M&M section.
@@
The noise reduction factor (NRF), defined as $\text{NRF}=~\frac{\langle\delta^{2}(N_{s}-N_{i})\rangle}{\langle N_{s}+N_{i}\rangle},$ is a well-established quantifier of non-classical correlations and shot-noise suppression, comparing the residual fluctuations in the difference of photon counts (numerator) to the shot-noise level expected for classical states (denominator) [7, 9, 29, 6, 3]. For classical light, the NRF satisfies $\text{NRF}\geq~1$, while quantum-correlated states can reach values in the sub-shot-noise regime, $0\leq\text{NRF}<1$.
In the case of two correlated regions of a detector collecting SPDC photons, the NRF can be expressed as $\text{NRF}=1-\eta_{0}\eta_{c}(\mathcal{D}),$ where $\eta_{0}$ is the single-channel detection efficiency, with the assumption of a balanced setup, $\eta_{s}=\eta_{i}=\eta_{0}$, and $\eta_{c}(\mathcal{D})$ represents the collection efficiency of correlated photon pairs [56, 54]. The quantity $\eta_{c}(\mathcal{D})$ is an increasing function of the ratio $\mathcal{D}=\frac{L_{\text{det}}}{M\cdot l_{\text{CFF}}},$ with $L_{\text{det}}$ the linear size of the detector’s integration area, $M$ the system magnification, and $l_{\text{CFF}}$ the correlation length in the crystal far-field. Further details about the estimation of $\eta_{0}$ and $\eta_{c}(\mathcal{D})$ are provided in the M&M section. When $\mathcal{D}\gg 1$, most of the correlated photon pairs are collected by two symmetric detection areas in the signal and idler beams, maximizing $\eta_{c}(\mathcal{D})\approx 1$; conversely, for $\mathcal{D}\lesssim 1$, only a fraction of the pairs are detected.
It is important to note that, while the spatial resolution $r_{\mathrm{CFF}}$ of the signal beam alone is inherently limited by the numerical aperture of the imaging system to $r_{\mathrm{CFF}}\approx 1.5\,\mu\mathrm{m}$ in our setup, noise subtraction in the intensity image is effective only on spatial scales larger than $l_{\mathrm{CFF}}$, that is for $\mathcal{D}\gtrsim 1$. This introduces a fundamental trade-off between spatial resolution and quantum advantage in SSNQI for amplitude objects, based on the estimator in Eq. (1). This dependency has been extensively characterized in previous works [56, 36, 54]. Under the assumptions of quasi-Poissonian photon number statistics in each arm and high object transmittance $\tau\approx 1$, the optimization factor $k^{(\tau)}_{\text{opt}}$ can be approximated as $k^{(\tau)}_{\text{opt}}(\mathcal{D})\approx 1-\text{NRF}=\eta_{0}\eta_{c}(\mathcal{D}),$ indicating the dependence on the resolution parameter $\mathcal{D}$. In practice, $k^{(\tau)}_{\text{opt}}$ is calibrated through a preliminary characterization of the experimental NRF as a function of $\mathcal{D}$ (see M&M).